A vinegar solution has a density of 1g/ml and contains 4.93% acetic by weight
How many moles of acetic acid are in 250 ml of this solution
1
Expert's answer
2019-12-03T08:24:44-0500

m(CH3COOH)=0,0493 × 250g = 12,325g

n(CH3COOH) = 12,325 ÷ 60,05 = 0,205 moles